# VisionLLaMA: A Unified LLaMA Interface for Vision Tasks

### v1

@@ 4.4 Object Detection on COCO
4.4.1 Supervised Training
We evaluate the performance of pyramid VisionLLaMA on the COCO objection detection task. Specifically, we use the Mask RCNN framework [26] and replace the backbone with pyramid VisionLLaMA, which is pre-trained for 300 epochs on the ImageNet-1K dataset as [43, 12]. Therefore, our model has the same number of parameters and FLOPs as Twins. Since our target is not to achieve a new state-of-the-art detector, this carefully controlled experiment is used to verify the validity of our method without loss of generality. Our implementation is based on the MMDetection framework [7] and the hyperparameter setting is provided in Section B.8. We report the result on standard 36 epochs (3$\times$) in Table 8. Under this carefully controlled setting, our model outperforms both Swin and Twins. Specifically, VisionLLaMA-B exceeds Swin-S by 1.5% box mAP and 1.0 mask mAP. Compared with the stronger baseline Twins-B, our method also has an advantage of 1.1% higher box mAP and 0.8% higher mask mAP.
Backbone FLOPs (G) Mask R-CNN 3$\times$ + MS
@@ 4.4.1 Supervised Training
Swin-S [43] 222 47.6 69.4 52.5 42.8 66.5 46.4
Twins-SVT-B[12] 224 48.0 69.5 52.7 43.0 66.8 46.6
Pyramid VisionLLaMA-B 224 49.1 70.5 54.0 43.8 67.4 47.0
Table 8: Object detection and instance segmentation performance on the COCO val2017 dataset using the Mask R-CNN framework. FLOPs are evaluated on an 800$\times$600 image. All the backbones are trained for 300 epochs on the ImageNet-1K dataset.
4.4.2 Self-Supervised Training
We apply VisionLLaMA based on the ViTDet framework[39], which utilizes plain vision transformers to achieve comparable performance as the pyramid counterpart. Specifically, we use the Mask RCNN detector and replace the vit-Base backbone (trained for 1600 epochs using MAE) with our VisionLLaMA-Base model, which is pre-trained for 800 epochs using MAE. The original ViTDet converges slowly and requires dedicated training strategies like longer training epochs (e.g. 100) to achieve optimal performance. During the training process, we find VisionLLaMA achieves similar performance after 30 epochs. Therefore, we directly utilize the standard 3x training strategy. We use AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We also use a layer-wise learning rate of 0.7 as [39]. The initial learning rate is 0.0001 and decayed by 0.1 at epochs 27 and 33. We use a weight decay of 0.1 and a global batch size of 64. The input image resolution is 1024$\times$1024. Therefore, our training cost is only 36% of the baseline. Unlike [39], we do not search for the optimal hyperparameter. The result is shown in Table 9 and VisionLLaMA outperforms ViT-B by 0.6% Box mAP and 0.8 % mask mAP.
[…]
@@ 4.4.2 Self-Supervised Training
Table 9: Object detection result on COCO 2017 dataset based on ViTDet[39]. sup: supervised training on ImageNet-1K
@@
5 Ablation Study and Discussion
5.1 Ablation Studies
Unless otherwise specified, we choose the ViT-Large model (160I 800E+224I 20E) to perform ablations because we observe that it generates small variance across multiple runs, where a performance gap of more than 0.2 suffices as a guide to choosing appropriate components.
case Acc
SwiGLU 84.6
FFN 84.6
Ratio Acc
25% 84.5
50% 84.5
100% 84.6
Base Acc
100 84.6
1000 84.6
10000 84.6
100000 84.4
Shared PE Acc
N 84.2
Y 84.6
Method Class Head Acc
VisionLLaMA-S Class Token 81.6
VisionLLaMA-S GAP 81.8
VisionLLaMA-B Class Token 83.6
VisionLLaMA-B GAP 83.6
VisionLLaMA-L Class Token 84.6
VisionLLaMA-L GAP 84.3
DeiT3-L [65] Class Token 84.5
${}^{\dagger}$ GAP 84.2
case Acc
Pyramid LLaMA-S 81.6
Pyramid LLaMA-S + learnable PE [64] 81.6
Pyramid LLaMA-S + PEG [13] 81.8
case Acc Train Speed
LayerNorm[2] 84.6 0.4971s
RMSNorm [80] 84.4 0.4874s
Table 11: Ablation experiments with plain transformer ViT-L/16 (DeiT3-L) on ImageNet-1K. We report the top-1 accuracy (%). If not specified, the default is: and the pre-training length is 800 epochs under an image resolution of 160$\times$160 and 20 epochs using 224$\times$224. Default settings are marked in gray. normal-$\dagger$: running the release code. All accuracies are top-1.
Ablation of FFN and SwiGLU. We replace FFN with SwiGLU and report the result in Table LABEL:tab:_ablaion_SwiGLU. We do not observe performance gaps, therefore, we utilize SwiGLU and avoid introducing extra modifications to the LLaMA architecture. This also motivates us to focus on the ablation of the self-attention block. As we apply multi-head self-attention, the remaining two differences become the normalization and positional encoding.
[…]
Next, we evaluate the role of positional encoding in two aspects, a static case using a fixed resolution and a dynamic case using variable resolutions. The former is common in the classification task while the latter is vital in downstream tasks such as segmentation and object detection.
Partial PE. We adjust the ratio of overall channels using RoPE to report the result in Table LABEL:tab:_partial_ratio, which shows good performance can be achieved if the ratio is set above a small threshold value. We do not observe significant differences across these settings. Therefore, we keep the default setting of [66] and do not follow [4, 30].
Frequency base. We change the base frequency and report the result in Table LABEL:tab:abalation_freq_base, which means the performance is robust to a large range of frequencies. As a result, we keep the default value of [66] to avoid extra special treatments for deployment.
@@ 5.1 Ablation Studies
Shared PE for each head. We find that sharing the same PE across different heads (the frequency varies from 1 to 10000 in each head) is better than independent ones (the frequency varies from 1 to 10000 across all channels). The result is shown in Table LABEL:tab:ablation_share_rpe.
@@
Feature abstraction strategy. We compare the two common feature extraction strategies: class token [22] and GAP [13] using the plain ‘large’ model and report the result in Table LABEL:tab:_ablation_feature_extraction. Using a class token is better than GAP, which is different from [13]. However, the training settings of the two cases are quite different. We also make an extra experiment using DeiT3-L to observe a similar performance gap of 0.3%. We further evaluate the performance of the ‘small’ and ‘base’ models. It’s interesting to see the opposite conclusions for the small model. We suspect that the higher drop-path rate used in [65] makes it difficult for the parameter-free abstraction such as GAP to fit in the purpose.
Positional encoding strategy. We also add other absolute position encoding strategies such as a learnable PE [64] and PEG [13] on pyramid VisionLLaMA-S. We use the ‘small’ model due to the existence of a strong baseline and report the result in Table LABEL:tab:_ablaion_pe. While the learnable PE does not boost performance, PEG slightly improves the baseline from 81.6% to 81.8%. However, we do not include PEG as a basic component regarding three aspects. Firstly, we try to keep the smallest modifications on LLaMA [66]. Secondly, our target is proposing a universal approach for various tasks like ViT [22]. For masked image frameworks like MAE [25], it is non-trivial to keep the reduced training cost of masked tokens if the backbone contains PEG. If we mask patches in the input like [75], it would greatly slow down the training speed. Moreover, containing masked patches in the encoder would incur a data distribution shift to the encoder, which severely hurts the performance of downstream tasks. In principle, we can apply sparse PEG under the MAE framework, but it will introduce the deployment-unfriendly operators. It remains an open problem whether sparse convolution contains enough positional information as its dense version [33, 13]. Thirdly, avoiding modality-bound designs paves the way for further studies that cover other modalities beyond text and vision.
Sensitivity to the input size. We further compare the performance on the enlarged and commonly used resolutions without training to report the result in Table 12. Here we use the pyramid transformer since it is more popular in downstream tasks than the plain counterpart. It is not surprising that 1D-RoPE severely suffers from the changed resolutions. NTK-Aware interpolation with $\alpha=2$ achieves similar performance as the 2D-RoPE [^1] Although we can apply the dynamic NTK-Aware to keep the performance at 224, it does not bring in boosted performance on larger resolutions., which is indeed NTK-Aware ($\alpha=1$). AS2DRoPE shows the best performance for larger resolution.
@@
Table 12: Top-1 accuracy on different resolutions of the pyramid small model. The models are trained on 224x224 and directly evaluated on other resolutions.
5.2 Discussion
We further investigate the underlying mechanisms behind our method’s superior performance over ViT in various tasks. As the ablation studies have indicated, our positional encoding strategy makes a big difference. In this section, we discuss the boosted convergence speed and attempt to theoretically rationalize the underlying mechanism.
Convergence speed. For image generation, we study the performance w.r.t the training steps. Specifically, we store the checkpoint at 100k, 200k, 300k, and 400k iterations to calculate the fidelity metrics. Since SDE is significantly slower than ODE, we opt to use the ODE sampler instead. The result of the strictly controlled experiment is listed in Table 10. It appears that VisionLLaMA converges much faster than ViT across all models. SiT-LLaMA with 300k training iterations even outperforms the baseline with 400k steps.
We also compare the convergence speed using the DeiT3-Large under the supervised training setting on ImageNet to show the top-1 validation accuracy during the 800 epochs in Figure 4. It also indicates that VisionLLaMA converges faster than DeiT3-L. We further compare the training loss across 800 epochs of the ViT-Base model under the MAE framework [25] and illustrate it in Figure 5. VisionLLaMA has lower training loss at the beginning and the trend is kept till the end.
Theoretical Reasoning. We dive into the mechanism of our positional encodings from the theoretical viewpoint. Without loss of generality, given an input embedding of dimension $d=4$, the query at location $(i,j)$ can be written as $q_{i,j}$. We use $k_{i,j}$ to represent the key vector at $(i,j)$ and $p_{i,j}$ to be the positional encoding using 2D sin-cos encoding[25, 46]. The inner dot product between $q_{i_{1},j_{1}}$ and $k_{i_{2},j_{2}}$ using this additive encoding can be written as,
$\begin{split}q_{i_{1},j_{1}}^{T}k_{i_{2},j_{2}}&=(q_{i_{1},j_{1}}+p_{i_{1},j_{1}})^{T}(k_{i_{2},j_{2}}+p_{i_{2},j_{2}})\\ &=q_{i_{1},j_{1}}^{T}k_{i_{2},j_{2}}+p_{i_{1},j_{1}}^{T}p_{i_{2},j_{2}}+q_{i_{1},j_{1}}^{T}p_{i_{2},j_{2}}+p_{i_{1},j_{1}}^{T}k_{i_{2},j_{2}}\\ &=q_{i_{1},j_{1}}^{T}k_{i_{2},j_{2}}+f(i_{1}-i_{2},j_{1}-j_{2})+M.\end{split}$ (5)
The first item is the inner dot product of contents. The second item reflects the positional effect in the form of $f(i_{1}-i_{2},j_{1}-j_{2})$, which plays a long-distance decaying effect. However, the third item $M=q_{i_{1},j_{1}}^{T}p_{i_{2},j_{2}}+p_{i_{1},j_{1}}^{T}k_{i_{2},j_{2}}$ means positions directly interacting with the content features, which slows down the learning process.
In contrast, the inner dot product using RoPE can be written as,
$\begin{split}(R_{i_{1},j_{1}}q_{i_{1},j_{1}})^{T}(R_{i_{2},j_{2}}k_{i_{2},j_{2}})&=q_{i_{1},j_{1}}^{T}R^{T}_{i_{1},j_{1}}R_{i_{2},j_{2}}k_{i_{2},j_{2}}\\ &=q_{i_{1},j_{1}}^{T}R_{i_{1}-i_{2},j_{1}-j_{2}}k_{i_{2},j_{2}}.\end{split}$ (6)
$R_{i_{1}-i_{2},j_{1}-j_{2}}$ contributes a larger absolute value if the positions of $q$ and $k$ are close, and a smaller value if opposite. This introduces certain localities as a prior bias, which resembles the function of a convolution. Moreover, $R_{i_{1}-i_{2},j_{1}-j_{2}}$ adjusts the dot product by the multiplication of a factor between 0 and 1, which is more flexible and faster than the addition of $f(i_{1}-i_{2},j_{1}-j_{2})$. We believe that this flexibility allows the transformer to leverage its model capacity effectively, learning a good representation without dedicating some of that capacity to introducing bias or separating position from content. In this way, VisionLLaMA not only converges faster but also has better final performance.
6 Conclusion
In a nutshell, we present VisionLLaMA to enjoy the benefits of the LLaMA architecture in the vision modality. It is trained either in supervised or self-supervised schemes to validate the power in a myriad of downstream vision tasks like image classification, detection, and segmentation. We particularly explore its image generation capacity under the diffusion framework DiT and SiT to confirm its potency. We conclude that VisionLLaMA has strong potential to serve as a new vision backbone to facilitate a large realm of downstream applications.
Acknowledgements:
Appendix A More Experiments
[…]
SiT-S/2 ${}^{\dagger}$ 6.06 33 400 0.0001 59.60 9.16 39.41 58.48 23.32
SiT-LLaMA-S/2 6.06 33 400 0.0001 54.62 8.81 41.69 61.16 26.07
SiT-B/2 ${}^{\dagger}$ 23.01 130 400 0.0001 36.90 6.70 51.00 64.10 39.78
SiT-LLaMA-B/2 23.01 130 400 0.0001 30.23 6.36 54.99 64.90 48.34
DiT-L/2 80.71 458 400 0.0001 23.3 - - - -
SiT-L/2 ${}^{\dagger}$ 80.71 458 400 0.0001 20.14 5.34 61.53 64.53 67.08
SiT-LLaMA-L/2 80.71 458 400 0.0001 14.91 5.16 64.97 64.30 82.23
SiT-XL/2 ${}^{\dagger}$ 118.64 675 400 0.0001 17.83 5.13 63.52 64.61 73.64
SiT-LLaMA-XL/2 118.64 675 400 0.0001 12.79 5.02 66.77 64.37 90.93
Table 13: Image generation comparisons using the SiT framework [46] All the models are trained using an image resolution of 256$\times$256 with a global batch size of 256. Metrics are calculated using the sampled 50k images without classifier-free guidance. IS: inception score. The FID is calculated by 250 steps ODE sampler because of the efficiency, which is a bit different from [46]. $\dagger$: reproduced result using the released code.
Appendix B Hyperparameters
B.1 Supervised Training of VisionLLaMA on ImageNet-1K

### v2

@@ 4.4 Object Detection on COCO
4.4.1 Supervised Training.
We evaluate the performance of pyramid VisionLLaMA on the COCO objection detection task. Specifically, we use the Mask RCNN framework [26] and replace the backbone with pyramid VisionLLaMA, which is pre-trained for 300 epochs on ImageNet-1K as [42, 12]. Since our target is not to achieve a new state-of-the-art detector, this carefully controlled experiment is used to verify the validity of our method. The hyperparameter setting is provided in Section 0.D.8(supp.). We report the result in Table 5. Our model outperforms both Swin and Twins. Specifically, VisionLLaMA-B exceeds Swin-S by 1.5% box mAP and 1.0 mask mAP. Compared with the stronger baseline Twins-B, ours also has an advantage of 1.1% higher box mAP and 0.8% higher mask mAP.
Table 5: Performance on the COCO2017 dataset using Mask R-CNN.
Backbone Setting Param (M) Mask R-CNN 3$\times$ + MS
@@
Swin-S [42] ImageNet1k 300E 69.1 47.6 69.4 52.5 42.8 66.5 46.4
Twins-SVT-B[12] ImageNet1k 300E 76.3 48.0 69.5 52.7 43.0 66.8 46.6
Pyramid VisionLLaMA-B ImageNet1k 300E 76.3 49.1 70.5 54.0 43.8 67.4 47.0
4.4.2 Self-Supervised Training.
We apply VisionLLaMA based on ViTDet [38], which utilizes plain transformers to achieve comparable performance as the pyramid counterpart. Specifically, we replace the vit-Base backbone (trained for 1600 epochs using MAE) with our VisionLLaMA-Base model, which is pre-trained for 800 epochs. The original ViTDet converges slowly and requires dedicated training strategies like longer training epochs (e.g., 100) to achieve good performance. During the training process, we find VisionLLaMA achieves similar performance after 30 epochs. Therefore, we directly utilize the standard 3$\times$ training strategy. Therefore, our training cost is only 36% of the baseline. Unlike [38], we do not search for the optimal hyperparameter. The result is shown in Table 6 and VisionLLaMA outperforms ViT-B by 0.6% Box mAP and 0.8% mask mAP.
Table 6: Object detection result on COCO 2017 dataset based on ViTDet[38].
Model Pretrained mAPBox mAPMask Epochs
@@
5 Ablation Study and Discussion
5.1 Ablation Studies
Unless otherwise specified, we choose the ViT-Large model (160I 800E+224I 20E) to perform ablations because we observe that it generates small variance across multiple runs, where a performance gap of more than 0.2 suffices as a guide to choosing appropriate components. We give ablations in Table 7 and more ablations are provided in Section 0.B (supp.).
Table 7: Ablation experiments with plain transformer ViT-L/16 (DeiT3-L) on ImageNet-1K. We report the top-1 accuracy (%). If not specified, the default is: and the pre-training length is 800 epochs under an image resolution of 160$\times$160 and 20 epochs using 224$\times$224. Default settings are marked in gray. $\dagger$: running the release code. All accuracies are top-1.
(a) SwiGLU/GELU
case Acc
SwiGLU 84.6
GELU 84.6
(b) Normalization.
case Acc Speed
LayerNorm[3] 84.6 0.4971s
RMSNorm [78] 84.4 0.4874s
(c) RoPE Ratio
Ratio Acc
25% 84.5
50% 84.5
100% 84.6
(d) RoPE base freq.
Base Acc
100 84.6
1000 84.6
10000 84.6
100000 84.4
(e) Shared PE across different heads. Shared PE for all heads is better.
Shared PE Acc
N 84.2
Y 84.6
(f) Feature extraction. The class token is better than GAP in the training setting of DeiT3[63].
Method Class Head Acc
VisionLLaMA-S Class Token 81.6
VisionLLaMA-S GAP 81.8
VisionLLaMA-B Class Token 83.6
VisionLLaMA-B GAP 83.6
VisionLLaMA-L Class Token 84.6
VisionLLaMA-L GAP 84.3
DeiT3-L [63] Class Token 84.5
DeiT3-L GAP 84.2
(g) PE comparison. Applying PEG [13] can further improve the performance. P-V-LLaMA: Pyramid VisionLLaMA, LPE: learnable PE
case Acc
P-LLaMA-S 81.6
P-V-LLaMA-S + LPE [62] 81.6
P-V-LLaMA-S + PEG [13] 81.8
Ablation of GELU and SwiGLU. We replace GELU with SwiGLU and report the result in Table 7(a). We do not observe performance gaps, therefore, we utilize SwiGLU and avoid introducing extra modifications to the LLaMA architecture. This also motivates us to focus on the ablation of the self-attention block. As we apply multi-head self-attention, the remaining two differences become the normalization and positional encoding.
[…]
Partial PE. We adjust the ratio of overall channels using RoPE to report the result in Table 7(c), which shows good performance can be achieved if the ratio is set above a small threshold value. We do not observe significant differences across these settings. Therefore, we keep the default setting of [64] and do not follow [5, 30].
Positional encoding strategy. We also add other absolute position encoding strategies such as a learnable PE [62] and PEG [13] on pyramid VisionLLaMA-S. We use the ‘small’ model due to the existence of a strong baseline and report the result in Table 7(g). While the learnable PE does not boost performance, PEG slightly improves the baseline from 81.6% to 81.8%. However, we do not include PEG as a basic component regarding three aspects. Firstly, we try to keep the smallest modifications on LLaMA [64]. Secondly, our target is proposing a universal approach for various tasks like ViT [22]. For masked image frameworks like MAE [25], it is non-trivial to keep the reduced training cost of masked tokens if the backbone contains PEG. If we mask patches in the input like [73], it would greatly slow down the training speed.
Sensitivity to the input size. We further compare the performance on the enlarged and commonly used resolutions without training to report the result in Table 12(supp.). Here we use the pyramid transformer since it is more popular in downstream tasks than the plain counterpart. It is not surprising that 1D-RoPE severely suffers from the changed resolutions. NTK-Aware interpolation with $\alpha=2$ achieves similar performance as the 2D-RoPE, which is indeed NTK-Aware ($\alpha=1$). AS2DRoPE shows the best performance for larger resolution.
5.2 Discussion
We further investigate the underlying mechanisms behind our method’s superior performance over ViT in various tasks. In this section, we discuss the boosted convergence speed and attempt to theoretically rationalize the mechanism.
Convergence speed. For image generation, we study the performance w.r.t the training steps. Specifically, we store the checkpoint at 100k, 200k, 300k, and 400k iterations to calculate the fidelity metrics. Since SDE is significantly slower than ODE, we opt to use the ODE sampler instead. The result of the strictly controlled experiment is listed in Table 10(supp.). It appears that VisionLLaMA converges much faster than ViT across all models. SiT-LLaMA with 300k training iterations even outperforms the baseline with 400k steps.
We also compare the convergence speed using the DeiT3-Large under the supervised training setting on ImageNet to show the top-1 validation accuracy during the 800 epochs in Figure 3(supp.). It also indicates that VisionLLaMA converges faster than DeiT3-L. We further compare the training loss across 800 epochs of the ViT-Base model under the MAE framework [25] and illustrate it in Figure 4(supp.). VisionLLaMA has lower training loss at the beginning and the trend is kept till the end.
Theoretical analysis. We dive into the mechanism of our positional encodings from the theoretical viewpoint. Without loss of generality, given an input embedding of dimension $d=4$, the query at location $(i,j)$ can be written as $q_{i,j}$. We use $k_{i,j}$ to represent the key vector at $(i,j)$ and $p_{i,j}$ to be the positional encoding using 2D sin-cos encoding[25, 45]. The inner dot product between $q_{i_{1},j_{1}}$ and $k_{i_{2},j_{2}}$ using this additive encoding can be written as,
$\begin{split}q_{i_{1},j_{1}}^{\!\top}k_{i_{2},j_{2}}&=(q_{i_{1},j_{1}}+p_{i_{1},j_{1}})^{\!\top}(k_{i_{2},j_{2}}+p_{i_{2},j_{2}})\\ &=q_{i_{1},j_{1}}^{\!\top}k_{i_{2},j_{2}}+p_{i_{1},j_{1}}^{\!\top}p_{i_{2},j_{2}}+q_{i_{1},j_{1}}^{\!\top}p_{i_{2},j_{2}}+p_{i_{1},j_{1}}^{\!\top}k_{i_{2},j_{2}}\\ &=q_{i_{1},j_{1}}^{\!\top}k_{i_{2},j_{2}}+f(i_{1}-i_{2},j_{1}-j_{2})+M.\end{split}$ (5)
The first item is the inner dot product of contents. The second item reflects the positional effect in the form of $f(i_{1}-i_{2},j_{1}-j_{2})$, which plays a long-distance decaying effect. However, the third item $M=q_{i_{1},j_{1}}^{\!\top}p_{i_{2},j_{2}}+p_{i_{1},j_{1}}^{\!\top}k_{i_{2},j_{2}}$ means positions directly interacting with the content features, which slows down the learning process.
In contrast, the inner dot product using RoPE can be written as,
$\begin{split}({\bf R}_{i_{1},j_{1}}q_{i_{1},j_{1}})^{\!\top}(\textbf{R}_{i_{2},j_{2}}k_{i_{2},j_{2}})&=q_{i_{1},j_{1}}^{\!\top}\textbf{R}^{\!\top}_{i_{1},j_{1}}\textbf{R}_{i_{2},j_{2}}k_{i_{2},j_{2}}\\ &=q_{i_{1},j_{1}}^{\!\top}\textbf{R}_{i_{1}-i_{2},j_{1}-j_{2}}k_{i_{2},j_{2}}.\end{split}$ (6)
$\textbf{R}_{i_{1}-i_{2},j_{1}-j_{2}}$ contributes a larger absolute value if the positions of $q$ and $k$ are close, and a smaller value if opposite. This introduces certain localities as a prior bias, which resembles the function of a convolution. Moreover, $\textbf{R}_{i_{1}-i_{2},j_{1}-j_{2}}$ adjusts the dot product by the multiplication of a factor between 0 and 1, which is more flexible and faster than the addition of $f(i_{1}-i_{2},j_{1}-j_{2})$. We believe that this flexibility allows the transformer to leverage its model capacity effectively, learning a good representation without dedicating some of that capacity to introducing bias or separating position from content. In this way, VisionLLaMA not only converges faster but also has better final performance.
6 Conclusion
In a nutshell, we present VisionLLaMA to enjoy the benefits of the LLaMA architecture in the vision modality. It is trained either in supervised or self-supervised schemes to validate the power in a myriad of downstream vision tasks including image classification, detection, and segmentation, among many others. We particularly explore its image generation capacity under the diffusion framework DiT and SiT to confirm its potency. We conclude that VisionLLaMA has a strong potential to serve as a new vision backbone to facilitate a large realm of downstream applications.
Acknowledgements
Appendix 0.A More Experiments
0.A.1 Image Generation
Table 8: Image generation of 256$\times$256 using SiT [45] without CFG. The FID is calculated by a 250-step SDE Euler sampler. $\dagger$: reproduced result using the released code.
[…]
(G) (M) (K)
SiT-S/2 6.06 33 400 0.0001 58.15 9.12 41.01 60.23 24.72
SiT-LLaMA-S/2 6.06 33 400 0.0001 53.90 8.78 42.98 60.36 26.74
SiT-B/2 23.01 130 400 0.0001 35.54 6.57 52.68 64.38 42.33
SiT-LLaMA-B/2 23.01 130 400 0.0001 29.53 6.32 56.07 64.07 50.13
DiT-L/2 80.71 458 400 0.0001 23.3 - - - -
SiT-L/2 80.71 458 400 0.0001 19.34 5.28 63.00 63.60 70.47
SiT-LLaMA-L/2 80.71 458 400 0.0001 14.32 5.17 66.39 63.64 86.85
SiT-XL/2 118.64 675 400 0.0001 16.98 5.07 65.12 64.10 77.06
SiT-LLaMA-XL/2 118.64 675 400 0.0001 12.20 5.03 67.86 63.08 95.28
Table 9: Image generation comparisons using the SiT framework [45] All the models are trained using an image resolution of 256$\times$256 with a global batch size of 256. Metrics are calculated using the sampled 50k images without classifier-free guidance. IS: inception score. The FID is calculated by a 250 steps ODE sampler because of the efficiency, which is a bit different from [45]. $\dagger$: reproduced result using the released code.
[…]
(G) (M) (K)
SiT-S/2 6.06 33 400 0.0001 59.60 9.16 39.41 58.48 23.32
SiT-LLaMA-S/2 6.06 33 400 0.0001 54.62 8.81 41.69 61.16 26.07
SiT-B/2 23.01 130 400 0.0001 36.90 6.70 51.00 64.10 39.78
SiT-LLaMA-B/2 23.01 130 400 0.0001 30.23 6.36 54.99 64.90 48.34
DiT-L/2 80.71 458 400 0.0001 23.3 - - - -
SiT-L/2 80.71 458 400 0.0001 20.14 5.34 61.53 64.53 67.08
SiT-LLaMA-L/2 80.71 458 400 0.0001 14.91 5.16 64.97 64.30 82.23
SiT-XL/2 118.64 675 400 0.0001 17.83 5.13 63.52 64.61 73.64
SiT-LLaMA-XL/2 118.64 675 400 0.0001 12.79 5.02 66.77 64.37 90.93
@@ 0.A.1 Image Generation
0.A.2 Image Classification
Model 160 224 256 288 512 768
DeiT3-Large [63] 83.1 84.5 84.7 84.6 82.1 76.5
VisionLLaMA-L 83.1 84.6 84.7 84.8 83.5 79.1
Appendix 0.B More Ablations
@@
Shared PE for each head. We find that sharing the same PE across different heads (the frequency varies from 1 to 10000 in each head) is better than independent ones (the frequency varies from 1 to 10000 across all channels). The result is shown in Table 7(e).
Frequency base. We change the base frequency and report the result in Table 7(d), which means the performance is robust to a large range of frequencies. As a result, we keep the default value of [64] to avoid extra special treatments for deployment.
@@
Feature abstraction strategy. We compare the two common feature extraction strategies: class token [22] and GAP [13] using the plain ‘large’ model and report the result in Table 7(f). Using a class token is better than GAP, which is different from [13]. However, the training settings of the two cases are quite different. We also make an extra experiment using DeiT3-L to observe a similar performance gap of 0.3%. We further evaluate the performance of the ‘small’ and ‘base’ models. It’s interesting to see the opposite conclusions for the small model. We suspect that the higher drop-path rate used in [63] makes it difficult for the parameter-free abstraction such as GAP to fit in the purpose.
Table 12: Top-1 accuracy on different resolutions of the pyramid small model. The models are trained on 224$\times$224 and directly evaluated on other resolutions.
@@
Appendix 0.C More Related Work
Appendix 0.D Hyperparameters
0.D.1 Supervised Training of VisionLLaMA on ImageNet-1K